# SpargeAttention: Accurate and Training-free Sparse Attention Accelerating Any Model Inference

### v7

@@ 1 Introduction
As sequence lengths in large models become longer, such as 45K-128K in video generation and language models (Yang et al., 2025b; Bao et al., 2024; Dubey et al., 2024; Zhang et al., 2025c), the time consuming of attention occupies a significant portion of inference latency in large models (Zhang et al., 2025d). Fortunately, the attention map $P=\mathrm{Softmax}(QK^{\top}/\sqrt{d})$ exhibits inherent sparsity, as the softmax operation often creates many values approaching zero (Deng et al., 2024). Sparse attention exploit such sparsity to accelerate attention by (1) constructing a “sparse mask”, which indicates the important non-zero entries of the attention map $P$ that should be computed, and (2) computing attention only for the parts corresponding to the sparse mask. There are three distinct categories of sparse attention methods based on how the sparse mask is generated. pattern-based method (Zhang et al., 2023; Xiao et al., 2024a; Fu et al., 2024; Zhu et al., 2024; Xiao et al., 2025, 2024b; Zhang et al., 2025h) relies on specific sparsity patterns based on empirical observations, dynamic sparse attention (Ribar et al., 2024; Singhania et al., 2024; Jiang et al., 2024; Lai et al., 2025; Gao et al., 2024; Xi et al., 2025; Yang et al., 2025a; Zhang et al., 2025f) computes the mask on-the-fly based on the inputs, and training-based method (Kitaev et al., 2020; Pagliardini et al., 2023; Zhang et al., 2025i) directly train models with native sparse attention.
@@
Limitation. (L1. Universality) Though existing sparse attention methods already demonstrate promising speedup on some tasks, their universality is still limited. Existing works are typically developed for specific tasks, such as language modeling, utilizing task-specific patterns such as sliding windows or attention sinks. However, the attention pattern varies significantly across tasks (see examples in Fig. 2), making these patterns hard to generalize. (L2. Usability) Moreover, it is difficult to implement both accurate and efficient sparse attention for any input. This is because accuracy demands precise prediction of the sparse regions in the attention map, while efficiency requires the overhead of this prediction to be minimal. However, current methods are difficult to effectively satisfy both of the requirements simultaneously. For example, MInference (Jiang et al., 2024) requires a large sequence length, such as 100K, to achieve a noticeable speedup.
@@
Our approach. In this work, we develop SpargeAttn, a training-free sparse attention that can be adopted universally on various tasks, including language modeling and text-to-image/video, and various sequence lengths. We propose three main techniques to improve the universality, accuracy, and efficiency. First, we propose a universal sparse mask prediction algorithm, which constructs the sparse mask by compressing each block of $Q$, $K$ to a single token. Importantly, we compress selectively based on the similarity of tokens within the block, so the algorithm can accurately predict sparse masks universally across tasks. Second, we propose a sparse online softmax algorithm at the GPU warp level, which further omits some PVPV products by leveraging the difference between global maximum values and local maximum values in online softmax. Third, we integrate this sparse approach into the 8-bit quantized SageAttention framework for further acceleration. To the best of our knowledge, SpargeAttn is the first sparse attention method that can actually accelerate across language, image, and video models without compromising accuracy.
@@ 2 Related Work
Depending on how the sparsity mask is constructed, sparse attention methods can be divided into three types: (1) Pattern required methods rely on some fixed patterns of the attention map, such as sliding windows or attention sinks (Xiao et al., 2024b). H2O (Zhang et al., 2023), InfLLM (Xiao et al., 2024a), and DUOAttention (Xiao et al., 2025) rely on sliding window pattern. SampleAttention (Zhu et al., 2024), MOA (Fu et al., 2024), and StreamingLLM (Xiao et al., 2024b) rely on sliding window and attention sink pattern. DitFastAttn (Yuan et al., 2024) relies on sliding window patterns and similarities between different attention maps. Moreover, DitFastAttn is restricted to simple diffusion transformers, showing incompatibility with language models and MMDiT models like Flux (Black Forest Labs, 2023), Stable Diffusion3 and 3.5 (Stability AI, 2023), and CogVideoX (Yang et al., 2025b). As the pattern varies across models, these methods may not universally work for different models. (2) Dynamic sparse methods dynamically construct the sparse mask based on the input without the need of preset patterns, and are thus potentially more universal. Existing works can be further categorized into channel compression and token compression. Channel compression methods include SparQAttn (Ribar et al., 2024) and LokiAttn (Singhania et al., 2024). They construct the mask by carrying full attention with reduced dimensionality. However, as the dimension is already small, e.g., 64, 128, in commonly used attention, the speedup potential might be limited. Token compression methods include MInference (Jiang et al., 2024) and FlexPrefill (Lai et al., 2025). They construct the mask by compressing each block of tokens to a single token and compute attention on this shorter sequence. However, this approximation is too aggressive: missing important blocks of $P$ is possible if they do not have a large attention score on the compressed sequence. SeerAttention (Gao et al., 2024) requires training of additional parameters for attention, which is expensive to use. Moreover, they are all designed for language models, and their applicability to other model types, such as diffusion models, remains uncertain. (3) Training-based methods modify the attention computation logic, requiring retraining the entire model, such as Reformer (Kitaev et al., 2020) and FastAttention (Pagliardini et al., 2023). These methods are much more expensive to use than training-free methods.
There are other ways to accelerate attention, such as optimizing the kernel implementation (Dao et al., 2022; Dao, 2024; Shah et al., 2024), quantization (Zhang et al., 2025d), distributing the workload (Liu et al., 2024a), and designing linear time attention (Wang et al., 2020; Choromanski et al., 2021; Yu et al., 2022; Katharopoulos et al., 2020). They are orthogonal to our approach.
@@ 3.1 Sparse FlashAttention
SpargeAttn adopts the tiling strategy of FlashAttention (Dao, 2024), and skip computing the blocks that are filtered out. Consider an attention operation $S=QK^{\top}/\sqrt{d},~P=\sigma(S),~O=PV$, where $\sigma(S)_{ij}=\exp(S_{ij})/\sum_{k}\exp(S_{ik})$ is the softmax operation. Let NN be the sequence length and $d$ be the dimensionality of each head; the matrices $Q$, $K$, and $V$ each have dimensions $N\times d$, while the matrix $S$ and $P$ is $N\times N$. FlashAttention proposes to tile $Q$, $K$, and $V$ from the token dimension into blocks $\{Q_{i}\},\{K_{i}\},\{V_{i}\}$ with block sizes $b_{q}$, $b_{k}$, $b_{k}$, respectively. Then, it uses online softmax (Milakov & Gimelshein, 2018) to progressively compute each block of $O$, i.e., $O_{i}$:
@@ 3.5 Combined with SageAttention
To further accelerate our implementation of sparse attention, we integrate our method into SageAttention (Zhang et al., 2025a, d, g, b, e), which proposes a quantized method for accelerating attention. Since quantization (Hu et al., 2025; Zhang et al., 2025j) operations and sparse operations are orthogonal, sparse computation can be directly applied to SageAttention. The complete algorithm is shown in Algorithm 1. Specifically, first, we need to add one judgment at the beginning of the inner loop of SageAttention (Line 10, Algorithm 1) to decide whether to skip the whole inner loop once. Second, we add another judgment before the updating of $O_{ij}$ in the inner loop of SageAttention (Line, in Algorithm 1) to decide whether to skip the computation of $\widetilde{P}_{ij}V_{j}$. Moreover, to minimize the attention map prediction overhead, we implement the prediction using CUDA and adopt some kernel fusion techniques.
@@ 4.1 Setup
Models. We validate the effectiveness of SpargeAttn across diverse representative models from language, image, and video generation. Specifically, we conduct experiments on Llama3.1 (8B) (Dubey et al., 2024) for text-to-text, CogvideoX (2B), Mochi (Team, 2024), and Open-Sora-Plan (Lin et al., 2024) for text-to-video, Flux (.1-dev) (Black Forest Labs, 2023) and Stable-Diffusion3.5 (large) (Stability AI, 2023) for text-to-image.
Datasets. The Text-to-text model is evaluated on four zero-shot tasks: WikiText (Merity et al., 2017) to assess the model’s prediction confidence, Longbench (Bai et al., 2024) and En.MC of InfiniteBench (Zhang et al., 2024) for a comprehensive assessment of long context understanding capabilities, and the Needle-in-a-Haystack task (Kamradt, 2023) to assess the model’s retrieval ability. Text-to-video models are evaluated using the open-sora (Zheng et al., 2024c) prompt sets. Text-to-image models are assessed on COCO annotations (Lin et al., 2014).
End-to-end metrics. For Llama3.1, we use perplexity (ppl.) (Jelinek et al., 1977) for WikiText, Longbench score (Bai et al., 2024), and retrival accuracy for the Needle-in-a-Haystack task (Kamradt, 2023). For text-to-video models, following Zhao et al. (2025b), we evaluate the quality of generated videos on five metrics: CLIPSIM and CLIP-Temp (CLIP-T) (Liu et al., 2024b) to measure the text-video alignment; VQA-a and VQA-t to assess the video aesthetic and technical quality, and Flow-score (FScore) for temporal consistency (Wu et al., 2023). For text-to-image models, generated images are compared with the images in the COCO dataset in three aspects: FID (Heusel et al., 2017) for fidelity evaluation, Clipscore (CLIP) (Hessel et al., 2021) for text-image alignment, and ImageReward (IR) (Xu et al., 2024) for human preference.
Speed and sparsity metric. We use TOPS (tera operations per second) to evaluate the speed of sparse attention methods. Specifically, TOPS = $O(attn)/t$, where $O(attn)$ represents the total number of operations in a standard attention computation, and $t$ is the latency from a given $(Q,K,V)$ to the output of attention. Note that this speed metric is completely fair. This is because the $O(attn)$ is fixed for a set of inputs, and then the speed is determined by $t$, which includes the time spent predicting the sparse region of the attention map. We define Sparsity as the proportion of the Matmul of $Q_{i}K_{j}^{\top}$ plus $\widetilde{P}_{ij}V_{j}$ that are skipped relative to the total number of $Q_{i}K_{j}^{\top}$ plus $\widetilde{P}_{ij}V_{j}$ in a full attention required.
@@
Baselines. Currently, sparse attention methods applicable across different model types are limited. We choose block-sparse MInference (Jiang et al., 2024) and FlexPrefill (Lai et al., 2025) as our baselines. To vary the sparsity of these baselines, we use 30% and 70% for MInference, and use $\gamma=0.95$ and $0.99$ for FlexPrefill according to their paper.
@@ A.1 Detailed Explain and results of permutation ablation
We use five distinct prompts and pre-searched hyperparameters with $l_{1}=0.05,l_{2}=0.06$ on both CogvideoX and Mochi models. The permutation are performed separately in attention operation for $Q,K,V$ after position embedding. To retain the original order of the input sequence, an inverse permutation is performed on the output of attention; for models using visual-language joint self-attention(e.g., CogvideoX), we only permute the visual tokens. When evaluating block self-similarity, we choose a block size of 128128 for query and 6464 for key, which aligns with our kernel implementation. The precision metric(L1) is evaluated using FlashAttention2 output as ground truth.
@@ A.4 Sparsity analysis over diffusion model
In this section, we analyze the sparsity patterns in CogvideoX across different dimensions: model layers, denoising timesteps, input samples, and attention heads. Figure 14 illustrates the layer-wise sparsity. Figure 15 demonstrates timestep-wise sparsity. Figure 16 highlights sample-wise sparsity. Figure 17 presents head-wise sparsity, illustrating the diversity in attention behavior across different heads. These analyses are helpful for the design of some diffusion algorithms (Zheng et al., 2023, 2024b, 2024a, 2025; Zhao et al., 2024, 2025a; Wang et al., 2024).

### v8

@@ 1 Introduction
As sequence lengths in large models become longer, such as 45K-128K in video generation and language models (yang2024cogvideox; bao2024vidu; llama31model; zhang2025sage), the time consuming of attention occupies a significant portion of inference latency in large models (zhangsurvey; 2024sageattention). Fortunately, the attention map $P=\mathrm{Softmax}(QK^{\top}/\sqrt{d})$ exhibits inherent sparsity, as the softmax operation often creates many values approaching zero (zhangsurvey). Sparse attention exploit such sparsity to accelerate attention by (1) constructing a “sparse mask”, which indicates the important non-zero entries of the attention map $P$ that should be computed, and (2) computing attention only for the parts corresponding to the sparse mask. There are three distinct categories of sparse attention methods based on how the sparse mask is generated. pattern-based method (zhang2023h2o; xiao2024infllm; moaattention; zhu2024sampleattention; xiao2024duoattention; xiao2023efficient; zhang2025fast) relies on specific sparsity patterns based on empirical observations, dynamic sparse attention (ribar2023sparq; singhania2024loki; jiang2407minference; FlexPrefill; gao2024seerattention; xi2025sparse; yang2025sparse; zhang2025spargeattn_wksp) computes the mask on-the-fly based on the inputs, and training-based method (kitaev2020reformer; pagliardini2023fast; zhang2025faster) directly train models with native sparse attention.
@@
Limitation. (L1. Universality) Though existing sparse attention methods already demonstrate promising speedup on some tasks, their universality is still limited. Existing works are typically developed for specific tasks, such as language modeling, utilizing task-specific patterns such as sliding windows or attention sinks. However, the attention pattern varies significantly across tasks (see examples in Fig. 2), making these patterns hard to generalize. (L2. Usability) Moreover, it is difficult to implement both accurate and efficient sparse attention for any input. This is because accuracy demands precise prediction of the sparse regions in the attention map, while efficiency requires the overhead of this prediction to be minimal. However, current methods are difficult to effectively satisfy both of the requirements simultaneously. For example, MInference (jiang2407minference) requires a large sequence length, such as 100K, to achieve a noticeable speedup.
@@
Our approach. In this work, we develop SpargeAttn, a training-free sparse attention that can be adopted universally on various tasks, including language modeling and text-to-image/video, and various sequence lengths. We propose three main techniques to improve the universality, accuracy, and efficiency. First, we propose a universal sparse mask prediction algorithm, which constructs the sparse mask by compressing each block of $Q$, $K$ to a single token. Importantly, we compress selectively based on the similarity of tokens within the block, so the algorithm can accurately predict sparse masks universally across tasks. Second, we propose a sparse online softmax algorithm at the GPU warp level, which further omits some $PV$ products by leveraging the difference between global maximum values and local maximum values in online softmax. Third, we integrate this sparse approach into the 8-bit quantized SageAttention framework for further acceleration. To the best of our knowledge, SpargeAttn is the first sparse attention method that can actually accelerate across language, image, and video models without compromising accuracy.
@@ 2 Related Work
Depending on how the sparsity mask is constructed, sparse attention methods can be divided into three types (zhangsurvey): (1) Pattern required methods rely on some fixed patterns of the attention map, such as sliding windows or attention sinks (xiao2023efficient). H2O (zhang2023h2o), InfLLM (xiao2024infllm), and DUOAttention (xiao2024duoattention) rely on sliding window pattern. SampleAttention (zhu2024sampleattention), MOA (moaattention), and StreamingLLM (xiao2023efficient) rely on sliding window and attention sink pattern. DitFastAttn (yuan2024ditfastattn) relies on sliding window patterns and similarities between different attention maps. Moreover, DitFastAttn is restricted to simple diffusion transformers, showing incompatibility with language models and MMDiT models like Flux (flux), Stable Diffusion3 and 3.5 (stable_diffusion_3_5), and CogVideoX (yang2024cogvideox). As the pattern varies across models, these methods may not universally work for different models. (2) Dynamic sparse methods dynamically construct the sparse mask based on the input without the need of preset patterns, and are thus potentially more universal. Existing works can be further categorized into channel compression and token compression. Channel compression methods include SparQAttn (ribar2023sparq) and LokiAttn (singhania2024loki). They construct the mask by carrying full attention with reduced dimensionality. However, as the dimension is already small, e.g., 64, 128, in commonly used attention, the speedup potential might be limited. Token compression methods include MInference (jiang2407minference) and FlexPrefill (FlexPrefill). They construct the mask by compressing each block of tokens to a single token and compute attention on this shorter sequence. However, this approximation is too aggressive: missing important blocks of $P$ is possible if they do not have a large attention score on the compressed sequence. SeerAttention (gao2024seerattention) requires training of additional parameters for attention, which is expensive to use. Moreover, they are all designed for language models, and their applicability to other model types, such as diffusion models, remains uncertain. (3) Training-based methods modify the attention computation logic, requiring retraining the entire model, such as Reformer (kitaev2020reformer) and FastAttention (pagliardini2023fast). These methods are much more expensive to use than training-free methods.
There are other ways to accelerate attention (zhangsurvey), such as optimizing the kernel implementation (dao2022flashattention; dao2023flashattention; shah2024flashattention), quantization (2024sageattention; zhang2024sageattention2; zhang2025sageattention2++; zhang2025sageattention3), distributing the workload (liu2023ringattentionblockwisetransformers), and designing linear time attention (wang2020linformer; choromanski2020rethinking; yu2022metaformer; katharopoulos2020transformers). They are orthogonal to our approach.
@@ 3.1 Sparse FlashAttention
SpargeAttn adopts the tiling strategy of FlashAttention (dao2023flashattention), and skip computing the blocks that are filtered out. Consider an attention operation $S=QK^{\top}/\sqrt{d},~P=\sigma(S),~O=PV$, where $\sigma(S)_{ij}=\exp(S_{ij})/\sum_{k}\exp(S_{ik})$ is the softmax operation. Let $N$ be the sequence length and $d$ be the dimensionality of each head; the matrices $Q$, $K$, and $V$ each have dimensions $N\times d$, while the matrix $S$ and $P$ is $N\times N$. FlashAttention proposes to tile $Q$, $K$, and $V$ from the token dimension into blocks $\{Q_{i}\},\{K_{i}\},\{V_{i}\}$ with block sizes $b_{q}$, $b_{k}$, $b_{k}$, respectively. Then, it uses online softmax (milakov2018online) to progressively compute each block of $O$, i.e., $O_{i}$:
@@ 3.5 Combined with SageAttention
To further accelerate our implementation of sparse attention, we integrate our method into SageAttention (zhang2024sageattention2; 2024sageattention; zhang2025sageattention2++; zhang2025sageattention2_wksp; zhang2025sageattention3), which proposes a quantized method for accelerating attention. Since quantization (hu2025quant; zhang2025int8train) operations and sparse operations are orthogonal, sparse computation can be directly applied to SageAttention. The complete algorithm is shown in Algorithm 1. Specifically, first, we need to add one judgment at the beginning of the inner loop of SageAttention (Line 10, Algorithm 1) to decide whether to skip the whole inner loop once. Second, we add another judgment before the updating of $O_{ij}$ in the inner loop of SageAttention (Line, in Algorithm 1) to decide whether to skip the computation of $\widetilde{P}_{ij}V_{j}$. Moreover, to minimize the attention map prediction overhead, we implement the prediction using CUDA and adopt some kernel fusion techniques.
@@ 4.1 Setup
Models. We validate the effectiveness of SpargeAttn across diverse representative models from language, image, and video generation. Specifically, we conduct experiments on Llama3.1 (8B) (llama31model) for text-to-text, CogvideoX (2B), Mochi (genmo2024mochi), and Open-Sora-Plan (lin2024open) for text-to-video, Flux (.1-dev) (flux) and Stable-Diffusion3.5 (large) (stable_diffusion_3_5) for text-to-image.
Datasets. The Text-to-text model is evaluated on four zero-shot tasks: WikiText (merity2022pointer) to assess the model’s prediction confidence, Longbench (bai2023longbench) and En.MC of InfiniteBench (zhang-etal-2024-bench) for a comprehensive assessment of long context understanding capabilities, and the Needle-in-a-Haystack task (LLMTest_NeedleInAHaystack) to assess the model’s retrieval ability. Text-to-video models are evaluated using the open-sora (opensora) prompt sets. Text-to-image models are assessed on COCO annotations (lin2014microsoft).
End-to-end metrics. For Llama3.1, we use perplexity (ppl.) (jelinek1977perplexity) for WikiText, Longbench score (bai2023longbench), and retrival accuracy for the Needle-in-a-Haystack task (LLMTest_NeedleInAHaystack). For text-to-video models, following zhao2024viditq, we evaluate the quality of generated videos on five metrics: CLIPSIM and CLIP-Temp (CLIP-T) (liu2024evalcrafter) to measure the text-video alignment; VQA-a and VQA-t to assess the video aesthetic and technical quality, and Flow-score (FScore) for temporal consistency (wu2023exploring). For text-to-image models, generated images are compared with the images in the COCO dataset in three aspects: FID (heusel2017gans) for fidelity evaluation, Clipscore (CLIP) (hessel2021clipscore) for text-image alignment, and ImageReward (IR) (xu2024imagereward) for human preference.
Speed and sparsity metric. We use inverse latency 1/t{1}/{t} to evaluate the speed of sparse attention methods. Specifically, 1/t{1}/{t} = $O(attn)/t$, where $O(attn)$ represents the total number of operations in a standard attention computation, and $t$ is the latency in seconds from a given $(Q,K,V)$ to the output of attention. Note that this speed metric is completely fair. This is because the $O(attn)$ is fixed for a set of inputs, and then the speed is determined by $t$, which includes the time spent predicting the sparse region of the attention map. We define Sparsity as the proportion of the Matmul of $Q_{i}K_{j}^{\top}$ plus $\widetilde{P}_{ij}V_{j}$ that are skipped relative to the total number of $Q_{i}K_{j}^{\top}$ plus $\widetilde{P}_{ij}V_{j}$ in a full attention required.
@@
Baselines. Currently, sparse attention methods applicable across different model types are limited. We choose block-sparse MInference (jiang2407minference) and FlexPrefill (FlexPrefill) as our baselines. To vary the sparsity of these baselines, we use 30% and 70% for MInference, and use $\gamma=0.95$ and $0.99$ for FlexPrefill according to their paper.
@@ A.1 Detailed Explain and results of permutation ablation
We use five distinct prompts and pre-searched hyperparameters with $l_{1}=0.05,l_{2}=0.06$ on both CogvideoX and Mochi models. The permutation are performed separately in attention operation for $Q,K,V$ after position embedding. To retain the original order of the input sequence, an inverse permutation is performed on the output of attention; for models using visual-language joint self-attention(e.g., CogvideoX), we only permute the visual tokens. When evaluating block self-similarity, we choose a block size of $128$ for query and $64$ for key, which aligns with our kernel implementation. The precision metric(L1) is evaluated using FlashAttention2 output as ground truth.
@@ A.4 Sparsity analysis over diffusion model
In this section, we analyze the sparsity patterns in CogvideoX across different dimensions: model layers, denoising timesteps, input samples, and attention heads. Figure 14 illustrates the layer-wise sparsity. Figure 15 demonstrates timestep-wise sparsity. Figure 16 highlights sample-wise sparsity. Figure 17 presents head-wise sparsity, illustrating the diversity in attention behavior across different heads. These analyses are helpful for the design of some diffusion algorithms (zheng2023dpm; zheng2024diffusion; zheng2024masked; zheng2025direct; zhao2024identifying; zhao2025riflex; wang2024framebridge).